OpenStudy (anonymous):

A product in 2008 was valued at $625.50 and the rate it is expected to change during the next five years is $42.70 increase per year. Use this information to write a linear equation giving the dollar value V of the product in terms of year t (t=8 represents 2008)

OpenStudy (anonymous):

v=t+42.50x let x= no of years

OpenStudy (anonymous):

I figured it out. The slope is m=42.70. The point it goes through is (8, 625.50). In the point-slope formula it is V-625.50=42.70(t-8) V=42.70t+283.90

OpenStudy (anonymous):

right...